# Galois Action and Localization in Number Fields

### v2

@@ 5. Localizations of Rings of Integers
It is sufficiently clear that conditions 1-3 from Definition 1.1 hold. Now, let $\mathfrak{P}]\in Cl(\mathcal{O}_{K}[\frac{1}{N(x)}])$ where $\mathfrak{P}$ is a prime ideal representative. If we write $\mathfrak{p}=\mathfrak{P}\cap\mathcal{O}_{K}$, we have $\left(\prod_{\sigma\in G}\sigma(\mathfrak{p})\right)\mathcal{O}_{K}[\frac{1}{N(x)}]=\prod_{\sigma\in G}\sigma(\mathfrak{P})$, so the norm property (4) is also inherited from our original group action. ∎
Now, any restrictions the structure of $G$ places on $Cl_{K}$ must also hold for $Cl(\mathcal{O}_{K}[\frac{1}{N(x)}])$ for any nonzero, nonunit $x$. In particular, if the set $\{[\mathfrak{p}_{1}],...,[\mathfrak{p}_{k}]\}$ is closed under the action of $G$, then we may apply our previous methods to $Cl_{K}/\left<[\mathfrak{p}_{1}],...,[\mathfrak{p}_{k}]\right$>. This is just another example of how misleading the quadratic case can be. As $\text{Gal}(K/\mathbb{Q})$ induces the subgroup $\{id,-id\}\leqslant\text{Aut}(Cl_{K})$ where $id$ is the identity and $id([I])=[I]^{-1}$, we see that any homomorphic image of $Cl_{K}$ can be realized by localizing at an integer—allowing significant restriction on $Cl_{K}$.
@@
This was proven for abelian number fields by Fröhlich ([9]) which is a notably more restrictive condition. For example, observe that while Theorem 3.2 would not preclude an extension of degree $K:\mathbb{Q}]=3^{r}$ from having $h_{k}=55$, Corollary 5.7 shows this is impossible. In fact, such an extension cannot have $h_{K}=5n$ for any $n\in\mathbb{N}$ not divisible by $5$.
@@ 6. Overrings of Dedekind Domains
Let $\mathcal{O}_{K}=\mathbb{Z}[\sqrt{-5}]$ and consider the overring $\mathcal{O}_{K}\left[\frac{1}{1+\sqrt{-5}}\right]$. Now, $\text{Gal}(K/\mathbb{Q})=\{\mathbf{1},\sigma\}$ where $\sigma$ represents complex conjugation. Note that $\sigma(R)$ fails to be an $R$-module as $\frac{1}{1+\sqrt{-5}}\cdot\sigma\left(\frac{1}{1+\sqrt{-5}}\right)=\frac{1}{1+\sqrt{-5}}\cdot\frac{1}{1-\sqrt{-5}}=\frac{1}{6}\notin\sigma(R)$.
The following is a classical result which characterizes all overrings of Dedekind domains with torsion class group. For details, the reader is encouraged to see [10].
Theorem 6.2.
Let $D$ be a Dedekind domain with torsion class group and $R$ an overring of $D$. Then
@@
From this, we can quickly derive the following theorem of which we will make great use.
@@
By Theorem 6.2, every overring of $D$ can be expressed in the form $\underset{\mathfrak{p}\in S}{\bigcap}D_{\mathfrak{p}}$ for some $S\subseteq\text{Spec}(D)$. Thus, it suffices to show that $\underset{\mathfrak{p}\in S_{1}}{\bigcap}D_{\mathfrak{p}}\neq\underset{\mathfrak{p}\in S_{2}}{\bigcap}D_{\mathfrak{p}}$ for subsets $S_{1},S_{2}\subseteq\text{Spec}(D)$ whenever $S_{1}\neq S_{2}$. Without loss of generality, assume $\mathfrak{q}\in S_{1}\backslash S_{2}.$ Recall that $Cl(D)$ is torsion, so $|[\mathfrak{q}]|:=n$ is finite. Thus, $\mathfrak{q}^{n}=(q)$ for some $q\in D$. Now, $\mathfrak{q}$ is the unique prime ideal containing $q$, so $\frac{1}{q}\in\underset{\mathfrak{p}\in S_{2}}{\bigcap}D_{\mathfrak{p}}\backslash\underset{\mathfrak{p}\in S_{1}}{\bigcap}D_{\mathfrak{p}}$. ∎
@@
By Theorem 6.3, we know that any adjacent domain $D\subseteq D^{\prime}$, there is a unique prime $\mathfrak{p}\in\text{Spec}(D)$ such that $\mathfrak{p}D^{\prime}=D^{\prime}$. What is more, any adjacent domain is a simple extension, and this $\mathfrak{p}$ determines the $\alpha$ such that $D^{\prime}=D[\alpha]$ as $\alpha=\frac{1}{a}$ where $aD=\mathfrak{p}^{n}$ where $|\mathfrak{p}|=n$. From these observations, we have the following Jordan-Hölder like result on overrings of Dedekind domains with finite class group.
@@
This theorem tells us that any ring of integers in particular is finitely many adjacent steps away from being a PID.
@@
Now, in the case of rings of integers, $R$ being a minimal PID over $\mathcal{O}_{K}$ does not imply the existence of such chain of adjacent domains. In fact, the following example implies that there always exists a PID over $\mathcal{O}_{K}$ failing to have this property except in the case that the ring of integers itself is a PID.
@@
Let $\mathcal{O}_{K}$ be a ring of integers with $Cl_{K}=\mathbb{Z}/2\mathbb{Z}$, and let $P$ be the set of all non-principal prime ideals of $\mathcal{O}_{K}$. Now, taking $S=\cup_{\mathfrak{p}\in P}\mathfrak{p}$, we have that $R_{S^{C}}=\{\frac{r}{a}\,|\,r\in R,a\notin S\}$ also has class group $\mathbb{Z}/2\mathbb{Z}$ by Theorem 1.5 of [3]. Now, take a (non-principal) prime $\mathfrak{q}\in\text{Spec}(R)$, and let $bR_{S}=\mathfrak{q}^{n}$ for some nn\ $\in\mathbb{N}$. Then, $R_{S}[\frac{1}{b}]$ is minimal over $\mathcal{O}_{K}$ with respect to being a PID, but there is no finite chain of adjacent domains between them by Theorem 6.3.
We now return to the particular case of rings of integers. For a more general treatment of overrings of Dedekind domains, the reader is referred to [3] and [4]. In the following theorem, we adopt the notation of [15] and denote the group of fractional ideals of a domain $R$ as $G(R)$.
@@
    $R$ is Galois-invariant. That is for all $\sigma\in G$, $\sigma(R)=R$.

### v4

@@ 4. Localization and the Class Group
It is sufficiently clear that conditions 1-3 from Definition 1.1 hold. Now, let $\mathfrak{P}]\in Cl(\mathcal{O}_{K}[\frac{1}{n}])$ where $\mathfrak{P}$ is a prime ideal representative. Note, such a $\mathfrak{P}$ must exist because each class of $Cl_{K}$ contains infinitely many primes and only finitely many intersect $\{\alpha^{n}\,|\,n\in\mathbb{N}$}. Thus, existence follows from the 1-1 correspondence between prime ideals of $\mathcal{O}_{K}[\frac{1}{\alpha}]$ and prime ideals of $\mathcal{O}_{K}$ disjoint from $S=\{\alpha^{n}\,|\,n\in\mathbb{N}\}$ given by $\mathfrak{p}\longleftrightarrow\mathfrak{p}\mathcal{O}_{K}[\frac{1}{\alpha}]$ for $\mathfrak{p}\in\text{Spec}(\mathcal{O}_{K})$.
If we write $\mathfrak{p}=\mathfrak{P}\cap\mathcal{O}_{K}$, we claim that $\left(\prod_{\sigma\in G}\sigma(\mathfrak{p})\right)\mathcal{O}_{K}[\frac{1}{n}]=\prod_{\sigma\in G}\sigma(\mathfrak{P})$, and thus the norm property (4) is also inherited from our original group action. We begin by noting that for any $\sigma\in G$, $\sigma(\mathcal{O}_{K})\subseteq\mathcal{O_{K}}$. Now, taking $\sigma^{-1}$ on both sides, we see that $\mathcal{O}_{K}\subseteq\sigma^{-1}(\mathcal{O}_{K})$ for any $\sigma\in G$, so in fact $\sigma(\mathcal{O}_{K})=\mathcal{O}_{K}$. Now, writing $G=\{\sigma_{i}\}_{i=1}^{n}$, let $x\in\left(\prod_{\sigma\in G}\sigma(\mathfrak{p})\right)\mathcal{O}_{K}[\frac{1}{n}]$ and write $x=\left(\prod_{i=1}^{n}\sigma_{i}(p_{i})\right)\cdot\frac{a}{n^{k}}$ where $p_{i}\in\mathfrak{p}$ and $a\in\mathcal{O}_{K}$. Let $b\in\mathcal{O}_{K}$ such that $\sigma(b)=a$. Then, $x=\sigma_{1}(\frac{p_{1}b}{n^{k}})\left(\prod_{i=2}^{n}\sigma_{i}(p_{i})\right)\in\prod_{\sigma\in G}\sigma(\mathfrak{P})$ as $\frac{p_{1}b}{n^{k}}\in\mathfrak{p}_{S}=\mathfrak{P}.$ For the other containment, let $y\in\prod_{\sigma\in G}\sigma(\mathfrak{P})$, and write $y=\prod_{i=1}^{n}\sigma_{i}(\frac{\bar{p}_{i}}{n^{r_{i}}})=\left(\prod_{i=1}^{n}\sigma_{i}(\bar{p}_{i})\right)\cdot\prod_{i=1}^{n}\frac{1}{n^{r_{i}}}\in\left(\prod_{\sigma\in G}\sigma(\mathfrak{p})\right)\mathcal{O}_{K}[\frac{1}{n}]$. ∎
The well-definedness of the action follows from the fact that $\sigma\left(\frac{1}{n}\right)=\frac{1}{n}\in\mathcal{O}_{K}[\frac{1}{n}]$ and thus $\sigma(\mathcal{O}_{K}[\frac{1}{n}])=\mathcal{O}_{K}[\frac{1}{n}]$ for any $\sigma\in\text{Gal}(K/\mathbb{Q})$. More on this in the next section. Now, any restrictions the structure of $G$ places on $Cl_{K}$ must also hold for $Cl(\mathcal{O}_{K}[\frac{1}{n}])$. In particular, on $Cl(\mathcal{O}_{K}[\frac{1}{N(\alpha)}]$ for any nonzero, nonunit $\alpha$. Thus, by Corollary 4.4, if the set $\{[\mathfrak{p}_{1}],...,[\mathfrak{p}_{k}]\}$ is closed under the action of $G$, then we may apply our previous methods to $Cl_{K}/\left<[\mathfrak{p}_{1}],...,[\mathfrak{p}_{k}]\right$>. This is just another example of how misleading the quadratic case can be. As $\text{Gal}(K/\mathbb{Q})$ induces the subgroup $\{id,-id\}\leqslant\text{Aut}(Cl_{K})$ where $id$ is the identity and $id([I])=[I]^{-1}$, we see that any homomorphic image of $Cl_{K}$ can be realized by localizing at an integer—allowing significant restriction on $Cl_{K}$.
@@
This was proven for abelian number fields by Fröhlich ([14]) which is a notably more restrictive condition. Now, observe that while Theorem 2.2 would not preclude an extension of degree $K:\mathbb{Q}]=3^{r}$ from having $h_{k}=55$, Corollary 4.8 shows this is impossible. In fact, such an extension cannot have $h_{K}=5n$ for any $n\in\mathbb{N}$ not divisible by $5$.
@@ 5. Overrings of Dedekind Domains
Let $\mathcal{O}_{K}=\mathbb{Z}[\sqrt{-5}]$ and consider the overring $\mathcal{O}_{K}\left[\frac{1}{1+\sqrt{-5}}\right]$. Now, $\text{Gal}(K/\mathbb{Q})=\{\textbf{1},\sigma\}$ where $\sigma$ represents complex conjugation. Note that $\sigma(R)$ fails to be an $R$-module as $\frac{1}{1+\sqrt{-5}}\cdot\sigma\left(\frac{1}{1+\sqrt{-5}}\right)=\frac{1}{1+\sqrt{-5}}\cdot\frac{1}{1-\sqrt{-5}}=\frac{1}{6}\notin R$.
The following is a well-known characterization of overrings of Dedekind domains which first appeared as Theorem 1 in [16].
Proposition 5.2.
Let $D$ be a Dedekind domain and $R$ an overring of $D$. Then
@@
The following result can be found in [17], and it follows quickly from the preceding proposition. We include a proof here as it is constructive and will inform later results and examples.
@@
By Proposition 5.2, every overring of $D$ can be expressed in the form $\underset{\mathfrak{p}\in S}{\bigcap}D_{\mathfrak{p}}$ for some $S\subseteq\text{Spec}(D)$. Thus, it suffices to show that $\underset{\mathfrak{p}\in S_{1}}{\bigcap}D_{\mathfrak{p}}\neq\underset{\mathfrak{p}\in S_{2}}{\bigcap}D_{\mathfrak{p}}$ for subsets $S_{1},S_{2}\subseteq\text{Spec}(D)$ whenever $S_{1}\neq S_{2}$. Without loss of generality, assume $\mathfrak{q}\in S_{1}\backslash S_{2}.$ Recall that $Cl(D)$ is torsion, so $|[\mathfrak{q}]|:=n$ is finite. Thus, $\mathfrak{q}^{n}=(q)$ for some $q\in D$. Now, $\mathfrak{q}$ is the unique prime ideal containing $q$, so $\frac{1}{q}\in\underset{\mathfrak{p}\in S_{2}}{\bigcap}D_{\mathfrak{p}}\backslash\underset{\mathfrak{p}\in S_{1}}{\bigcap}D_{\mathfrak{p}}$. ∎
@@
By the correspondence in Theorem 5.3, we see that for any adjacent ring $R$ over D, there is a unique $\mathfrak{p}\in\text{Spec}(D)$ such that $\mathfrak{p}R=R$. What is more, recalling that any adjacent extension simple, $\mathfrak{p}$ uniquely determines the irreducible $a\in D$ (up to a unit) such that $R=D[\frac{1}{a}]$. In particular, we have $(a)=\mathfrak{p}^{n}$ where $n=|[\mathfrak{p}]|$. Inspired by this uniqueness, we present the following Jordan–Hölder like result on sequences of adjacent overring.
@@
We will now shift our focus back to the particular case of rings of integers. For a more general treatment of overrings of Dedekind domains and their class groups, the reader is encouraged to see [9] and [17]. Theorem 5.5 tells us that any ring of integers in particular is finitely many adjacent steps away from being a PID. In this setting, this is equivalent to asking how far the ring is from being a UFD—a question considered more broadly in [2].
@@
Now, even in the case of rings of integers, $R$ being a minimal PID over $\mathcal{O}_{K}$ does not imply the existence of such chain of adjacent domains. In fact, the following example implies that there always exists a PID over $\mathcal{O}_{K}$ failing to have this property except in the case that the ring of integers itself is a PID.
@@
Let $\mathcal{O}_{K}$ be a ring of integers with $Cl_{K}=\mathbb{Z}/2\mathbb{Z}$, and let $P$ be the set of all non-principal prime ideals of $\mathcal{O}_{K}$. Now, taking $S=\cup_{\mathfrak{p}\in P}\mathfrak{p}$, we have that $R_{S}=\{\frac{r}{a}\,|\,r\in R,a\notin S\}$ also has class group $\mathbb{Z}/2\mathbb{Z}$ by Theorem 1.5 of [7]. Note, this is equivalent to localizing at the multiplicatively closed set generated by the prime elements of $\mathcal{O}_{K}$. Now, take a (non-principal) prime $\mathfrak{q}\in\text{Spec}(R)$, and let $aR_{S}=\mathfrak{q}^{n}$ for some nn\ $\in\mathbb{N}$. Then, $R_{S}[\frac{1}{a}]$ is minimal over $\mathcal{O}_{K}$ with respect to being a PID, but there is no finite chain of adjacent domains between them by Theorem 5.3.
We now return to the particular case of rings of integers. For a more general treatment of overrings of Dedekind domains, the reader is referred to [7] and [8]. In the following theorem, we adopt the notation of [24] and denote the group of fractional ideals of a Dedekind domain $R$ as $G(R)$.
@@
    $R$ is Galois-invariant. That is, for all $\sigma\in G$, $\sigma(R)=R$.